40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 7 09:00:26 2022, Paco, Update, General, Scheduled power outage recovery 
    Reply  Mon Feb 7 15:32:48 2022, Yehonathan, Update, General, Y End laser controller 
    Reply  Mon Feb 7 16:53:02 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220208_001646282.jpgPXL_20220208_001657871.jpg
       Reply  Wed Feb 9 11:56:24 2022, Anchal, Update, General, Bringing back CDS Screenshot_2022-02-09_12-11-33.png
          Reply  Wed Feb 9 13:55:05 2022, Koji, Update, General, Bringing back CDS 
       Reply  Wed Feb 9 16:43:35 2022, Paco, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
          Reply  Thu Feb 10 15:41:00 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
             Reply  Thu Feb 10 17:57:48 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) PXL_20220211_021509819.jpg
                Reply  Thu Feb 10 19:03:23 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                   Reply  Thu Feb 10 19:46:37 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                      Reply  Thu Feb 10 21:10:43 2022, Koji, Update, General, Video Mux setting reset Screenshot_2022-02-10_21-11-21.pngScreenshot_2022-02-10_21-11-54.png
             Reply  Fri Feb 11 16:09:11 2022, Anchal, Update, General, Scheduled power outage recovery - Input power increased 
                Reply  Mon Feb 14 18:31:50 2022, Paco, Update, General, Scheduled power outage recovery - IMC recovery progress 
                   Reply  Tue Feb 15 19:32:50 2022, Koji, Update, General, Scheduled power outage recovery - IMC recovery progress PXL_20220216_001731377.jpgScreen_Shot_2022-02-15_at_16.18.16.pngPXL_20220216_001727465.jpgScreen_Shot_2022-02-15_at_16.22.16.pngPXL_20220216_002229572.jpg
                      Reply  Tue Feb 15 19:40:02 2022, Koji, Update, General, IMC locking 
                         Reply  Wed Feb 16 15:19:41 2022, Anchal, Update, General, Reconfigured MC reflection path for low power 
                            Reply  Wed Feb 23 15:08:57 2022, Anchal, Update, General, Removed extra beamsplitter in MC WFS path 
                               Reply  Thu Feb 24 14:32:57 2022, Anchal, Update, General, MC RFPD DCMON channel got stuck to 0 
                                  Reply  Thu Feb 24 19:26:32 2022, Anchal, Update, General, IMC Locking 
                                     Reply  Sun Feb 27 00:37:00 2022, Koji, Update, General, IMC Locking Recovery PXL_20220226_093809056.jpgPXL_20220226_093854857.jpgPXL_20220226_100859871.jpgScreenshot_2022-02-26_01-56-31.pngScreenshot_2022-02-26_01-56-47.png
                                        Reply  Sun Feb 27 01:12:46 2022, Koji, Update, General, IMC manual alignment procedure PXL_20220226_100859871.jpg
    Reply  Fri Feb 11 11:17:00 2022, Anchal, Update, General, Scheduled power outage recovery 
       Reply  Mon Feb 14 21:03:25 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220215_025325118.jpg
Message ID: 16673     Entry time: Tue Feb 15 19:40:02 2022     In reply to: 16672     Reply to this: 16674
Author: Koji 
Type: Update 
Category: General 
Subject: IMC locking 

IMC is locking now. There was nothing wrong: just a careful alignment + proper gain adj

=== Primary Alignment ===

- I used WFS error signals as the indicator of the PDH error signals. Checked C1:IOO-WFS1_(I/Q)n_ERR and ended up C1:IOO-WFS1_I4_ERR as it showed the largest PDH error PP.

- Then used MC2 and MC3 to align the IMC by maximizing the PDH error and the MC trans (C1:IOO-MC_TRANS_SUM_ERR)

=== Locking procedure ===

Note that the MC REFL path is still configured for the full power input

- (Only at the beginning) Run scripts/MC/mcdown for initialization / Run scripts/MC/MC2tickleOFF just in case

- Enable IOO-MC-SW1 (MC SERVO switch right after "IN1 Gain (dB)").
- Disable 40:4000 boost
- Increase VCO Gain from -15 to 0
- Jiggle IN1 Gain from low to +31 until the lock is achieved

- As soon as the lock is acquired, enable 40:4000
- Increase VCO Gain to +10
- Turn up "SUPER BOOST" from 0 to 3

=== Lock loss procedure ===

Note that the MC REFL path is still configured for the full power input

- Disable IOO-MC-SW1
- Disable 40:4000 boost
- Reduce VCO Gain 0
- Turn down "SUPER BOOST" to 0

- Then jiggle IN1 Gain again to lock the IMC

=== MC2 spot ===

- It was obvious that the MC2F spot was not on the center of the optic.
- I tried to move the spot on the camera as much as possible, but this did not make the trans beam to the center of the MC end QPD
- I had the impression that the trans beam started to be clipped when the beam was moved towards the end QPD,

We need to reestablish the reasonable/consistent MC2 spot on the mirror, the MC end optics, and the QPD.
We will need to use MC2 dithering and A2L coupling to determine the center of the mirror

But as long as the transmission is maximized, the transmitted beam thru MC1 and MC3 follows the input beam. So we can continue the vent work

The current maximized transmission was ~1300. MC1 refl CCD view was largely off -> The camera path was adjusted.

=== MC2 alignment note ===

During the alignment, I noticed a sudden change of the MC2 alignment. There might be some hysteresis in the MC2 suspension. If you are locking the IMC and noticed significant misalignment, the first thing to try is to touch MC2 alignment.

ELOG V3.1.3-